# Resilient Distributed Optimization for Multi-Agent Cyberphysical Systems

### v2

@@ 7 Numerical Results
The trust values are generated as follows. Let $\mathbf{E}[\alpha_{ij}(t)]=0.55$ if $j\in\mathcal{N}_{i}\cap\mathcal{L}$, and $\mathbf{E}[\alpha_{ij}(t)]=0.45$ if $j\in\mathcal{N}_{i}\cap\mathcal{M}$. The random variable $\alpha_{ij}(t)$ is uniformly distributed on the interval $\left[\mathbf{E}[\alpha_{ij}(t)]-\frac{\ell}{2},\mathbf{E}[\alpha_{ij}(t)]+\frac{\ell}{2}\right]$, for every $i\in\mathcal{L}$ and $j\in\mathcal{L}\cup\mathcal{M}$. We consider the values $\ell$: $0.6,0.8$, in both scenarios $|E_{\mathcal{L}}|=|E_{\mathcal{M}}|=0.05$, however, the variance of the trust values when $\ell=0.8$ are higher. We remark that the legitimate agents are ignorant regarding the values $\mathbf{E}[\alpha_{ij}]$ and $\ell$. Due to the stochasticity of the trust values, we average the numerical results across $100$ system realizations. Furthermore, we use the following stepsize $\gamma(t)=\frac{2}{t+2}\cdot\mathds{1}_{\{t\geq 0\}}$.
@@
The W-MSR algorithm [11] for robust distributed optimization, which we use for comparison in the supplementary material for the one-dimensional case, is only valid for one-dimensional data values. Thus, as a benchmark we compare our results to the multi-dimensional extension of the W-MSR algorithm proposed in [63]. Following the notations in [11], we denote by $F$ the maximal number of highest values and lowest values that each legitimate agent discards, overall a legitimate agent may ignore no more than $2F$ values for $d=1$. We note that in this case the number of tolerated malicious inputs for the W-MSR algorithm is reduced by a factor of $d$, i.e., the dimension of the data values. Additionally, we use the nominal case with no malicious agents (5) as another benchmark to evaluate the performance of Algo. 1.
In our numerical results, we use the following auxiliary vector $(\tilde{b}_{i})_{i=1}^{15}=(115.7,163.3,-81.7,127.2,-63.7,58.4,-3.1,$ 62.9,54.5,144.9,121.1,9.3,2.6,124.5,131)62.9,54.5,144.9,-121.1,9.3,-2.6,-124.5,131)62.9 , 54.5 , 144.9 , - 121.1 , 9.3 , - 2.6 , - 124.5 , 131 ). Note that setting $d=1$, $a_{i}=1$, and $\lambda=0$ results in a one-dimensional constrained consensus problem. We consider this special case in Section III in the supplementary material.
@@
Figs. 3 and 3 capture the average value of the squared distance of each legitimate agent from the optimal point $x_{\mathcal{L}}^{\star}$, i.e., the average value of $\overline{e}(t)$, for each time $t$. The plots presented in Figs. 3 and 3 for our higher dimensional setup are consistent with the one-dimensional setup which is captured in Figs. 1-4 in the supplementary material. We can see that Algo. 1 performs well and mitigates the harmful effect of malicious inputs even in higher dimensions. This is in contrast to the multi-dimensional W-MSR algorithm [63] which is more vulnerable to malicious attacks as the dimension of the data values, i.e., $d$, increases.
In particular, Algo. 1 provides resilience to malicious activity and can tolerate even $30=2|\mathcal{L}|$ malicious agents, as evident in Figs. 3 and 3. Furthermore, we can see from Figs. 3 and 3 that Algo. 1 is robust to undesirably small values of $|E_{\mathcal{L}}|$ and $|E_{\mathcal{M}}|$ where the trust observations are less informative. Figs. 3 and 3 demonstrate how our algorithm eventually recovers the global optimum as predicted by theory regardless of the value of the observation window $T_{0}$. Thus, the value of $T_{0}$ mostly dictates the rate of recovery of the global optimum. Particularly, Figs. 3 and 3 show that the variance of the trust values has more impact on $T_{0}$ values that are smaller than $100$. This occurs since the higher variance of the trust values increases the misclassification errors. Since the probability of these errors decreases with $T_{0}$, they are less impactful when $T_{0}$ is $100$ or higher. Finally, we refer the reader to Section III in the supplementary material for additional numerical results.
@@ 8 Conclusions
where $(a)$ follows from the definition of $T_{f}$ in Corollary 1 which implies that if $T_{f}=k$ there must be a misclassification error in the legitimacy of agents at time $k-1$, $(b)$ follows from the union bound, and (c) follows from Lemma 1. Now, since by definition, $p_{c}(-1)\geq 1$ we can conclude (11).
From the definition of $T_{f}$ in Corollary 1 $\Pr(T_{f}>k-1)=1$ for $k=0$. Additionally, by the union bound, for all $k\geq 1$
@@
$\leq\sum_{t=k-1}^{\infty}\sum_{i\in\mathcal{L}}|\mathcal{N}_{i}\cap\mathcal{L}|\exp(-2tE_{\mathcal{L}}^{2})$
$\qquad+\sum_{t=k{\color[rgb]{0,0,1}-1}}^{\infty}\sum_{i\in\mathcal{L}}|\mathcal{N}_{i}\cap\mathcal{M}|\exp(-2tE_{\mathcal{M}}^{2})$
@@
$=\sum_{k=0}^{t-1}\Pr(T_{f}=k)\mathbf{E}\left[\|x_{\mathcal{L}}(t,T_{f}=k)-x_{\mathcal{L}}^{\star}\|^{r}\>\big{|}\>T_{f}=k\right]$
@@
$\quad+\Pr(T_{f}>t-1)\left(2\eta\right)^{r}.$
@@
$\leq$ $\sum_{k=0}^{t/2}\Pr(T_{f}=k)\left[\frac{4\overline{h}(t-M_{k})|\mathcal{L}|}{\mu(t-M_{k})(t-M_{k}+1)}\right]^{\frac{r}{2}}$
$+\sum_{k=t/2+1}^{t-1}\Pr(T_{f}=k)\cdot\min\left\{\left[\frac{4\overline{h}(t-M_{k})|\mathcal{L}|}{\mu(t-M_{k})(t-M_{k}+1)}\right]^{\frac{r}{2}},(2\eta)^{r}\right\}.$
Now, since $t\geq 2T_{0}$ and $\frac{\overline{h}(t)}{t(t+1)}$ is a decreasing function we have that
$\sum_{k=0}^{t/2}\Pr(T_{f}=k)\left[\frac{4\overline{h}(t-M_{k})|\mathcal{L}|}{\mu(t-M_{k})(t-M_{k}+1)}\right]^{\frac{r}{2}}$
$\leq\left[\frac{4\overline{h}(t/2)|\mathcal{L}|}{\mu(t/2)(t/2+1)}\right]^{\frac{r}{2}}\sum_{k=1}^{t/2}\Pr(T_{f}=k)\leq\left[\frac{4\overline{h}(t/2)|\mathcal{L}|}{\mu(t/2)(t/2+1)}\right]^{\frac{r}{2}},$
@@
$\sum_{k=t/2+1}^{t-1}\Pr(T_{f}=k)\cdot\min\left\{\left[\frac{4\overline{h}(t-M_{k})|\mathcal{L}|}{\mu(t-M_{k})(t-M_{k}+1)}\right]^{\frac{r}{2}},(2\eta)^{r}\right\}$
@@
Additionally,
@@
where $(a)$ follows from the triangle inequality, $(b)$ follows from (11), $(c)$ follows from the double stochasticity of $W(t)$, and $(d)$ follows from (11).
Thus,
@@
then $\mathbf{E}[\|\Delta_{i}(t)\|]\leq\delta(t)$ for all $i\in\mathcal{L}$.
It follows from (3)
@@
We conclude the proof by utilizing (33), (34). ∎

### v3

@@ 7 Numerical Results
The trust values are generated as follows. Let $\mathbf{E}[\alpha_{ij}(t)]=0.55$ if $j\in\mathcal{N}_{i}\cap\mathcal{L}$, and $\mathbf{E}[\alpha_{ij}(t)]=0.45$ if $j\in\mathcal{N}_{i}\cap\mathcal{M}$. The random variable $\alpha_{ij}(t)$ is uniformly distributed on the interval $\left[\mathbf{E}[\alpha_{ij}(t)]-\frac{\ell}{2},\mathbf{E}[\alpha_{ij}(t)]+\frac{\ell}{2}\right]$, for every $i\in\mathcal{L}$ and $j\in\mathcal{L}\cup\mathcal{M}$. We consider the values $\ell$: $0.6,0.8$, in both scenarios $|E_{\mathcal{L}}|=|E_{\mathcal{M}}|=0.05$, however, the variance of the trust values when $\ell=0.8$ are higher. We remark that the legitimate agents are ignorant regarding the values $\mathbf{E}[\alpha_{ij}]$ and $\ell$. Due to the stochasticity of the trust values, we average the numerical results across $100$ system realizations. Furthermore, we use the following stepsize $\gamma(t)=\frac{2}{(1-\epsilon)(t+2)}\cdot\mathds{1}_{\{t\geq 0\}},\>\>\epsilon=0.1$.
@@
The W-MSR algorithm [11] for robust distributed optimization, which we use for comparison in the supplementary material [63] for the one-dimensional case, is only valid for one-dimensional data values. Thus, as a benchmark we compare our results to the multi-dimensional extension of the W-MSR algorithm proposed in [64]. Following the notations in [11], we denote by $F$ the maximal number of highest values and lowest values that each legitimate agent discards, overall a legitimate agent may ignore no more than $2F$ values for $d=1$. We note that in this case the number of tolerated malicious inputs for the W-MSR algorithm is reduced by a factor of $d$, i.e., the dimension of the data values. Additionally, we use the nominal case with no malicious agents (5) as another benchmark to evaluate the performance of Algo. 1.
In our numerical results, we use the following auxiliary vector $(\tilde{b}_{i})_{i=1}^{15}=(115.7,163.3,-81.7,127.2,-63.7,58.4,-3.1,$ 62.9,54.5,144.9,121.1,9.3,2.6,124.5,131)62.9,54.5,144.9,-121.1,9.3,-2.6,-124.5,131)62.9 , 54.5 , 144.9 , - 121.1 , 9.3 , - 2.6 , - 124.5 , 131 ). Note that setting $d=1$, $a_{i}=1$, and $\lambda=0$ results in a one-dimensional constrained consensus problem. We consider this special case in Section III in the supplementary material [63].
@@
Figs. 3 and 3 capture the average value of the squared distance of each legitimate agent from the optimal point $x_{\mathcal{L}}^{\star}$, i.e., the average value of $\overline{e}(t)$, for each time $t$. The plots presented in Figs. 3 and 3 for our higher dimensional setup are consistent with the one-dimensional setup which is captured in Figs. 1-4 in the supplementary material [63]. We can see that Algo. 1 performs well and mitigates the harmful effect of malicious inputs even in higher dimensions. This is in contrast to the multi-dimensional W-MSR algorithm [64] which is more vulnerable to malicious attacks as the dimension of the data values, i.e., $d$, increases.
In particular, Algo. 1 provides resilience to malicious activity and can tolerate even $30=2|\mathcal{L}|$ malicious agents, as evident in Figs. 3 and 3. Furthermore, we can see from Figs. 3 and 3 that Algo. 1 is robust to undesirably small values of $|E_{\mathcal{L}}|$ and $|E_{\mathcal{M}}|$ where the trust observations are less informative. Figs. 3 and 3 demonstrate how our algorithm eventually recovers the global optimum as predicted by theory regardless of the value of the observation window $T_{0}$. Thus, the value of $T_{0}$ mostly dictates the rate of recovery of the global optimum. Particularly, Figs. 3 and 3 show that the variance of the trust values has more impact on $T_{0}$ values that are smaller than $100$. This occurs since the higher variance of the trust values increases the misclassification errors. Since the probability of these errors decreases with $T_{0}$, they are less impactful when $T_{0}$ is $100$ or higher. Finally, we refer the reader to Section III in the supplementary material [63] for additional numerical results.
@@ 8 Conclusions
where $(a)$ follows from the definition of $T_{f}$ in Corollary 1 which implies that if $T_{f}=k$ there must be a misclassification error in the legitimacy of agents at time $k-1$, $(b)$ follows from the union bound, and (c) follows from Lemma 1.
Now, since by definition, $p_{c}(-1)\geq 1$ we can conclude (11).
From the definition of $T_{f}$ in Corollary 1 $\Pr(T_{f}>k-1)=1$ for $k=0$. Additionally, by the union bound, for all $k\geq 1$,
@@
$\leq\sum_{t=k-1}^{\infty}\>\sum_{i\in\mathcal{L}}|\mathcal{N}_{i}\cap\mathcal{L}|\exp(-2tE_{\mathcal{L}}^{2})$
$\qquad+\sum_{t=k-1}^{\infty}\>\sum_{i\in\mathcal{L}}|\mathcal{N}_{i}\cap\mathcal{M}|\exp(-2tE_{\mathcal{M}}^{2})$
@@
$=\sum_{k=0}^{t-1}\Pr(T_{f}=k)\mathbf{E}\left[\|z_{i}(t-M_{k})-x_{\mathcal{L}}^{\star}\|^{r}\>\big{|}\>T_{f}=k\right]$
@@
$\quad+\Pr(T_{f}>t-1)\left(2\eta\right)^{r},$
where $z_{i}(0)=x_{i}(M_{k})$ for $i\in\mathcal{L}$.
@@
$\leq$ $\sum_{k=0}^{t/2}\Pr(T_{f}=k)\left[\frac{4\overline{h}_{\epsilon}(t-M_{k})|\mathcal{L}|}{\mu_{\epsilon}(t-M_{k})(t-M_{k}+1)}\right]^{\frac{r}{2}}$
$+\sum_{k=t/2+1}^{t-1}\Pr(T_{f}=k)\cdot\min\left\{\left[\frac{4\overline{h}_{\epsilon}(t-M_{k})|\mathcal{L}|}{\mu_{\epsilon}(t-M_{k})(t-M_{k}+1)}\right]^{\frac{r}{2}},(2\eta)^{r}\right\}.$
Now, since $t\geq 2T_{0}$, $\mu_{\epsilon}>0$ and $\frac{\overline{h}_{\epsilon}(t)}{t(t+1)}$ is a decreasing function we have that
$\sum_{k=0}^{t/2}\Pr(T_{f}=k)\left[\frac{4\overline{h}_{\epsilon}(t-M_{k})|\mathcal{L}|}{\mu_{\epsilon}(t-M_{k})(t-M_{k}+1)}\right]^{\frac{r}{2}}$
$\leq\left[\frac{4\overline{h}_{\epsilon}(t/2)|\mathcal{L}|}{\mu_{\epsilon}(t/2)(t/2+1)}\right]^{\frac{r}{2}}\sum_{k=0}^{t/2}\Pr(T_{f}=k)\leq\left[\frac{4\overline{h}_{\epsilon}(t/2)|\mathcal{L}|}{\mu_{\epsilon}(t/2)(t/2+1)}\right]^{\frac{r}{2}}.$
@@
$\sum_{k=t/2+1}^{t-1}\Pr(T_{f}=k)\cdot\min\left\{\left[\frac{4\overline{h}_{\epsilon}(t-M_{k})|\mathcal{L}|}{\mu_{\epsilon}(t-M_{k})(t-M_{k}+1)}\right]^{\frac{r}{2}},(2\eta)^{r}\right\}$
@@
Additionally, by the singular value decomposition of the doubly-stochastic matrix $W(t)$ we have that
@@
where $(a)$ follows from the triangle inequality, $(b)$ follows from (11), $(c)$ follows from the double stochasticity of $W(t)$, and $(d)$ follows from (11). Thus,
@@
it follows that $\mathbf{E}[\|\Delta_{i}(t)\|]\leq\delta(t)$ for all $i\in\mathcal{L}$.
From (3) we have
@@
We conclude the proof by utilizing (33) and (34). ∎